# Pessimism Principle Can Be Effective: Towards a Framework for Zero-Shot Transfer Reinforcement Learning

### v1

@@ Abstract
Keywords Zero-Shot  $\cdot$ Transfer Learning  $\cdot$ Distributed Learning  $\cdot$ Reinforcement Learning
@@ 1 Introduction
Reinforcement learning (RL) aims to learn a policy that optimizes an agent’s performance when interacting with its environment. RL has achieved remarkable success in areas such as game playing [1, 2, 3], robotics [4, 5] or natural language processing [6, 7, 8]. These successes, however, often depend on access to vast amounts of data, which are critical for enabling agents to understand their environments and develop effective policies. In domains with limited data—due to high exploration costs or other constraints—RL performance typically degrades. A promising solution is transfer learning (TL), which leverages more data from related environments (source domains) to train policies that can be transferred to target environments. We focus on zero-shot transfer, where no target domain data is available prior to deployment, and multi-domain transfer, where multiple source domains are utilized.
Despite its potential, TL is often hindered by the Sim-to-Real Gap [9, 10], which refers to performance degradation when transferring learned policies from source to target domains. TL typically relies on constructing a proxy to approximate target domain performance and transferring the proxy’s optimal policy. While successful when source domains closely resemble the target domain (e.g., high-fidelity simulations), significant divergence between the source and target domains—caused by modeling errors, non-stationarity, or perturbations—leads to inaccurate proxies and degraded transfer performance.
@@
These challenges are critical in practical applications, where ineffective policies may have severe consequences, such as mechanical damage in robotics or traffic accidents in autonomous driving. Existing TL methods, such as domain randomization (DR) or imitation learning, lack systematic guarantees for transferred performance, as they fail to establish a robust connection between the proxy and target domain outcomes. To address these limitations, and inspired by the effectiveness of pessimism principle in various RL settings, particularly in scenarios where conservativeness is advantageous, such as offline RL [11, 12], we propose a novel framework that incorporates the pessimism principle in proxy construction for TL. By using a pessimistic estimation of target domain performance, our approach ensures that transferred policies achieve satisfactory outcomes while avoiding severe consequences from overly optimistic actions. Our framework also mitigates negative transfer in multi-domain settings. We rigorously characterize the effectiveness of this approach and design concrete algorithms with convergence guarantees. Our contributions are summarized as follows.
Introducing pessimism principle in transfer learning.
We introduce the pessimism principle to transfer learning, enabling the transfer of conservative yet effective policies to the target domain. We characterize the framework’s effectiveness, showing that the target domain’s optimality gap depends on the degree of pessimism used in proxy construction. This motivates the design of proxies with minimal but sufficient pessimism, ensuring robust lower bounds on transferred performance and avoiding the severe consequences common in other TL methods.
Construction of pessimistic proxy and design of concrete algorithms.
Leveraging the inherent pessimism in robust RL frameworks, we construct pessimistic proxies by incorporating prior knowledge of domain similarities, such as upper bounds on the distance between source and target domains. We develop several proxy types, demonstrate their effectiveness, and propose a distributed algorithm for multi-domain transfer. This algorithm ensures convergence and yields conservative but effective policies across diverse source domains.
Design of proxy and algorithm to avoid negative transfer.
In multi-domain transfer, some source domains may significantly diverge from the target, leading to overly pessimistic policies if treated equally. We show that the pessimism principle inherently mitigates negative transfer by improving the proxy’s value without overly optimistic assumptions. Building on this insight, we propose a refined proxy that avoids negative transfer and design a convergent algorithm to ensure effective policy transfer in such settings.
@@ 2 Preliminaries and Problem Formulation
which exists and can be restricted to the set of deterministic policies [13]. The optimal value functions are $V_{P}^{\ast}\triangleq\max_{\pi}V_{P}^{\pi}=V^{\pi^{*}}_{P}$ and $Q_{P}^{\ast}\triangleq\max_{\pi}Q_{P}^{\pi}=Q^{\pi^{*}}_{P}$
Robust Markov Decision Process (RMDP)
Robust RL aims to optimize the performance under the worst case, when the environment is uncertain, and it is generally formulated as a robust MDP. The transition kernel in a robust MDP is not fixed but lies in an uncertainty set $\mathcal{P}$. In this paper, we focus on $(s,a)$-rectangular uncertainty sets [14, 15]. Given a nominal kernel $P_{0}$, the uncertainty set $\mathcal{P}$ is defined as
@@ 3 Related Work
Imitation Learning and Policy Distillation
Imitation learning (IL) and policy distillation are extensively explored TL techniques that enable an agent to learn from demonstrations or from simpler policies. In IL, an agent mimics expert behavior through supervised learning, often using behavioral cloning or inverse RL [16, 17, 18, 19]. Policy distillation transfers knowledge from a teacher policy to a student policy by matching the teacher’s action distribution [20, 21, 22, 23, 24]. The success of these approaches, however, rely on the assumption that the source domain’s expert or teacher policy is highly aligned with the target domain, which may not always be the case.
Domain Randomization
Domain randomization (DR) is widely used to reduce the Sim-to-Real Gap, by training agents on randomized versions of the simulated domain [25, 26, 27, 28, 29]. By introducing degrees of variability in domain properties such as physics, or object appearances, the agent aims to optimize the expected performance under a uniform distribution over these environments. DR is expected to enhance the robustness and generalizability of the learned policy so that it performs well when deployed in the target domain. Although DR has proven effective in many tasks [30, 31, 32, 33], it has limitations when the domain shift is large or the target environment is significantly different from the simulated one, in which case the randomized environments do not provide enough information for the target domain. Furthermore, DR generally lacks theoretical guarantees and justifications for the performance in the target domain, except under some strict assumptions about the structures of the underlying MDPs [34, 35].
Multi-task Learning
Multi-task learning (MTL) is another method that aims to leverage knowledge from multiple related tasks to improve learning efficiency and generalization. Specifically, MTL aims to optimize a finite set of tasks simultaneously or to obtain a Pareto stationary point [36, 37, 38, 39]. When the target domain shares common properties with these source tasks, this information will be maintained during the optimization of all source tasks and transferred. However, similar to DR, MTL also suffers from insufficient theoretical justification, requiring assumptions on the common representations of tasks. Moreover, MTL cannot avoid negative transfer, where some source tasks can be relatively distinct from the target task, thus slow down transfer learning. Also, solving an MTL problem can be difficult, due to the differences among the tasks known as the gradient conflict [40].
@@ 4.1 Major Barriers in Prior TL Methods
As discussed, most existing TL approaches optimize a proximal objective function $f(\pi)$ and deploy its optimal policy in the target domain. However, there are limited guarantees regarding the relationship between the proxy $f(\pi)$ and the target domain value function $V^{\pi}_{P_{0}}$, except under some strong assumptions (e.g., [34]). Consequently, it is unclear whether the optimal policy for $f(\pi)$ will also perform well under $P_{0}$, leaving uncertainty in its performance. Furthermore, when multiple source domains are available, existing methods struggle to distinguish domains that are more dissimilar to the target, often treating all domains equally. This inability to prioritize relevant sources can lead to negative transfer, where the knowledge transferred from additional source domains hurt the performance of the target domain instead of improving it.
@@ 4.3 A Robust RL Based Pessimism Principle
Robust RL is inherently pessimistic when tackling environment uncertainty. Specifically, when the environment of interest lies within the defined uncertainty set, the robust value function naturally provides a conservative lower bound on performance. This property inspires our studies of robust RL in transfer learning, as constructing an uncertainty set that includes the target domain is not challenging, and thus we can obtain a conservative proxy. Generally, the learner has some prior knowledge about the domain similarities to motivate transfer learning, and such similarities are usually captured in terms of the source and target domain kernels. For instance, a common characterization is the distance/divergence between the source and target kernels, e.g., [41]. Based on this information, an uncertainty set can be constructed to include the target domain kernel, and result in its robust value function as a conservative proxy.
@@
It can be shown that all of these proxies are conservative estimations of $V^{\pi}_{P_{0}}$, however, none of these proxies can be efficiently solved in a distributed setting. Specifically, (robust) DR is generally intractable and is approximated by a proximal DR objective [42]. Meanwhile, solving other proxies may require sharing local source domain models or data, which is often impractical or undesirable in a distributed setting. Therefore, we propose two conservative proxies based on robust RL, and show that they can be efficiently optimized under the distributed setting.
@@ 5.1 Averaged Operator Based proxy
Notice that according to the proof of Proposition 5.4, the result holds provided that $\sigma_{\mathcal{P}_{s}^{a}}-\frac{1}{K}\sum_{k=1}^{K}\sigma_{(\mathcal{P})_{s}^{a}}(V)\leq 0$. In Addition, for $l_{p}$-norm, the support function has a duality: $\sigma(V)=\max_{\alpha}\{PV_{\alpha}+f(\Gamma,V_{\alpha})\}$ for some function $f$ [43]. Thus $\sigma_{\mathcal{P}_{s}^{a}}-\frac{1}{K}\sum_{k=1}^{K}\sigma_{(\mathcal{P})_{s}^{a}}(V)=\max_{\alpha}\{\bar{P}V_{\alpha}+f(\Gamma,V_{\alpha})\}-\frac{1}{K}\sum_{k=1}^{K}\max_{\alpha}\{P_{k}V_{\alpha}+f(\Gamma,V_{\alpha})\}\leq 0$, and the similar result still holds.
@@ 5.2 Efficient and Distributed Algorithm Design
Recall $\mathbf{T}_{k}$ being the optimal robust Bellman operator for the $k$-th source domain as (4). Define the averaged optimal Bellman operator as
@@
In practice, the exact source domains may be unknown, but extensive samples from them are available, as the source domains are more free to explore. We hence assume that each local source domain can obtain an unbiased estimate of its robust Bellman operator: for any vector $Q$, an unbiased estimate $\hat{\mathbf{T}}_{k}Q$ is available to each local agent. This is a mild assumption, as there are different ways to construct an unbiased estimator of the robust Bellman operator, e.g., [44, 45, 46, 47]. We also note that due to the linear structure of $\mathbf{T}_{\rm AO}$, its unbiased estimation can be directly constructed as the average of all local estimations as $\hat{\mathbf{T}}_{\rm AO}=\frac{\sum^{K}_{k=1}\hat{\mathbf{T}}_{k}}{K}.$
On the other hand, to further avoid extensive communication, the frequency of aggregation can be reduced. Specifically, we update the local estimation of each source domain individually for $E$ steps, and then aggregate all the domains once. Such a learning scheme is also used in federated or distributed learning, e.g., [42, 48, 49].
@@
As the result shows, our algorithm enjoys a partial linear speedup, implying the efficiency of our algorithm for multi-domain transfer. The second term measures the trade-off between the convergence rate and communication cost, which is also observed in federated learning (FL) with heterogeneous environments [48]. We note that the convergence rate of our stochastic algorithm still enjoys a partial linear speedup, which matches the minimax optimal robust up to polylog factors [50]. It hence illustrates the efficiency of our algorithm, in terms of convergence rate and communication cost.
@@ 6 Avoiding Negative Transfer: the Minimal Pessimism Principle
While our averaged operator-based method provides a conservative yet effective proxy for target domain performance, it remains vulnerable to a key challenge in multi-domain transfer: negative transfer. Similar to existing TL methods, the averaged operator-based proxy treats all source domains equally and cannot identify or down-weight the less relevant ones. Inspired by recent advances in personalized FL [51, 52, 53, 54], the global learner should assign higher weights to source domains that are more similar to the target domain. However, identifying these similar domains without additional information remains challenging.
@@
(1). Both ${\mathbf{T}}^{\pi}_{{\rm MP}}$ and ${\mathbf{T}}_{{\rm MP}}$ are $\gamma$-contractions, and have unique fixed points, $Q^{\pi}_{{\rm MP}}$ and $Q_{{\rm MP}}$, respectively.
(2). For any policy $\pi$, let $V^{\pi}_{{\rm MP}}(s)=\sum_{a\in\mathcal{A}}\pi(a|s)Q^{\pi}_{{\rm MP}}(s,a)$, then $V^{\pi}_{\mathcal{P}_{k}}\leq V^{\pi}_{{\rm MP}}\leq V^{\pi}_{P_{0}}$. Moreover, $V^{\pi}_{\rm AO}\leq V^{\pi}_{{\rm MP}}$.
(3). Let $\pi_{\ast}(s)=\arg\max_{a\in\mathcal{A}}Q_{{\rm MP}}(s,a)$, then $Q_{{\rm MP}}=Q^{\pi_{\ast}}_{{\rm MP}}$ and $\pi_{\ast}=\arg\max_{\pi}Q_{{\rm MP}}^{\pi}$.
@@
To address this issue, it is essential to design an alternative maximal aggregation scheme that ensures convergence despite the presence of estimation errors. We propose an aggregation strategy to overcome this challenge. Specifically, we adopt the multi-level Monte-Carlo method [55, 56, 57] to construct an operator $\hat{\mathbf{M}}_{\rm MLMC}$ of the maximal aggregation based on estimated local operators. We defer the detailed construction to Appendix E.1.
@@
The MLMC module for MDTL-Max will increase the sample/computation complexity, which can be viewed as the price of improving effectiveness. Nevertheless, it can be reduced along two potential directions: One is to control the level number through techniques like threshold-MLMC [58]. Although it results in a biased estimation, the bias can be controlled and hence still implies convergence (see Section B.4). Another one is to relax the uncertainty set constraint. For example, for total variation, the relaxation results in the solution $P_{0}V-\Gamma\text{Span}(V)\leq\sigma(V)$ [59], which is much easier to compute and remains conservative. Reducing the aggregation frequency is an alternative way to reduce the computational cost, but also introduces a trade-off in the convergence rate, as we will show in the following results.

### v2

@@ Abstract
Zero-Shot, Transfer Learning, Distributed Learning, Reinforcement Learning
@@ 1 Introduction
Reinforcement learning (RL) aims to learn a policy that optimizes an agent’s performance when interacting with its environment. RL has achieved remarkable success in areas such as game playing (Silver et al., 2016, 2017; Mnih et al., 2013), robotics (Nguyen & La, 2019; Liu et al., 2021) or natural language processing (Sharma & Kaushik, 2017; Uc-Cetina et al., 2023; Kirk et al., 2023a). These successes, however, often depend on access to vast amounts of data, which are critical for enabling agents to understand their environments and develop effective policies. In domains with limited data—due to high exploration costs or other constraints—RL performance typically degrades. A promising solution is transfer learning (TL), which leverages more data from related environments (source domains) to train policies that can be transferred to target environments. We focus on zero-shot transfer, where no target domain data is available prior to deployment, and multi-domain transfer, where multiple source domains are utilized.
Despite its potential, TL is often hindered by the Sim-to-Real Gap (Salvato et al., 2021; Zhao et al., 2020), which refers to performance degradation when transferring learned policies from source to target domains. TL typically relies on constructing a proxy to approximate target domain performance and transferring the proxy’s optimal policy. While successful when source domains closely resemble the target domain (e.g., high-fidelity simulations), significant divergence between the source and target domains—caused by modeling errors, non-stationarity, or perturbations—leads to inaccurate proxies and degraded transfer performance.
@@
These challenges are critical in practical applications, where ineffective policies may have severe consequences, such as mechanical damage in robotics or traffic accidents in autonomous driving. Existing TL methods, such as domain randomization (DR) or imitation learning, lack systematic guarantees for transferred performance, as they fail to establish a robust connection between the proxy and target domain outcomes. To address these limitations, and inspired by the effectiveness of pessimism principle in various RL settings, particularly in scenarios where conservativeness is advantageous, such as offline RL (Jin et al., 2020; Shi et al., 2022), we propose a novel framework that incorporates the pessimism principle in proxy construction for TL. By using a pessimistic estimation of target domain performance, our approach ensures that transferred policies achieve satisfactory outcomes while avoiding severe consequences from overly optimistic actions. Our framework also mitigates negative transfer in multi-domain settings. We rigorously characterize the effectiveness of this approach and design concrete algorithms with convergence guarantees. Our contributions are summarized as follows.
Introducing pessimism principle in transfer learning. We introduce the pessimism principle to transfer learning, enabling the transfer of conservative yet effective policies to the target domain. We characterize the framework’s effectiveness, showing that the target domain’s optimality gap depends on the degree of pessimism used in proxy construction. This motivates the design of proxies with minimal but sufficient pessimism, ensuring robust lower bounds on transferred performance and avoiding the severe consequences common in other TL methods.
Construction of pessimistic proxy and design of concrete algorithms. Leveraging the inherent pessimism in robust RL frameworks, we construct pessimistic proxies by incorporating prior knowledge of domain similarities, such as upper bounds on the distance between source and target domains. We develop several proxy types, demonstrate their effectiveness, and propose a distributed algorithm for multi-domain transfer. This algorithm ensures convergence and yields conservative but effective policies across diverse source domains.
Design of proxy and algorithm to avoid negative transfer. In multi-domain transfer, some source domains may significantly diverge from the target, leading to overly pessimistic policies if treated equally. We show that the pessimism principle inherently mitigates negative transfer by improving the proxy’s value without overly optimistic assumptions. Building on this insight, we propose a refined proxy that avoids negative transfer and design a convergent algorithm to ensure effective policy transfer in such settings.
@@ 2 Preliminaries and Problem Formulation
which exists and can be restricted to the set of deterministic policies (Puterman, 2014). The optimal value functions are $V_{P}^{\ast}\triangleq\max_{\pi}V_{P}^{\pi}=V^{\pi^{*}}_{P}$ and $Q_{P}^{\ast}\triangleq\max_{\pi}Q_{P}^{\pi}=Q^{\pi^{*}}_{P}$
Robust Markov Decision Process (RMDP). Robust RL aims to optimize the performance under the worst case, when the environment is uncertain, and it is generally formulated as a robust MDP. The transition kernel in a robust MDP is not fixed but lies in an uncertainty set $\mathcal{P}$. In this paper, we focus on $(s,a)$-rectangular uncertainty sets (Nilim & El Ghaoui, 2004; Iyengar, 2005). Given a nominal kernel $P_{0}$, the uncertainty set $\mathcal{P}$ is defined as
@@ 3 Related Work
Domain Randomization: Domain randomization (DR) is widely used to reduce the Sim-to-Real Gap, by training agents on randomized versions of the simulated domain (Tobin et al., 2017; Vuong et al., 2019; Mehta et al., 2020; Tiboni et al., 2023; Tommasi, 2023). By introducing degrees of variability in domain properties such as physics, or object appearances, the agent aims to optimize the expected performance under a uniform distribution over these environments. DR is expected to enhance the robustness and generalizability of the learned policy so that it performs well when deployed in the target domain. Although DR has proven effective in many tasks (Shakerimov et al., 2023; Niu et al., 2021; Slaoui et al., 2019; Jiang et al., 2023), it has limitations when the domain shift is large or the target environment is significantly different from the simulated one, in which case the randomized environments do not provide enough information for the target domain. Furthermore, DR generally lacks theoretical guarantees and justifications for the performance in the target domain, except under some strict assumptions about the structures of the underlying MDPs (Chen et al., 2021; Jiang et al., 2022).
Multi-task Learning: Multi-task learning (MTL) is another method that aims to leverage knowledge from multiple related tasks to improve learning efficiency and generalization. Specifically, MTL aims to optimize a finite set of tasks simultaneously or to obtain a Pareto stationary point (Yang et al., 2020; Wilson et al., 2007; Vithayathil Varghese & Mahmoud, 2020; Teh et al., 2017). When the target domain shares common properties with these source tasks, this information will be maintained during the optimization of all source tasks and transferred. However, similar to DR, MTL also suffers from insufficient theoretical justification, requiring assumptions on the common representations of tasks. Moreover, MTL cannot avoid negative transfer, where some source tasks can be relatively distinct from the target task, thus slow down transfer learning. Also, solving an MTL problem can be difficult, due to the differences among the tasks known as the gradient conflict (Wang et al., 2024e).
Representation Learning for Multitask RL: Representation learning aims to extract the shared structure across multiple tasks, and utilize such a structure to reduce sample complexity in downstream tasks with similar features (Teh et al., 2017; Sodhani et al., 2021; Arulkumaran et al., 2022)Cheng et al. (2022); Agarwal et al. (2023); Sam et al. (2024) investigated multitask representation learning in an online RL setting, where the agent interacts with multiple source tasks to extract a common latent structure. Ishfaq et al. (2024) proposed Multitask Offline Representation Learning (MORL), which adopt the principle of pessimism and learn a shared representation from offline datasets across multiple tasks modeled by low-rank MDPs. However, these methods rely on the assumption that all tasks share a low-rank latent structure and that a common representation can be learned and reused, which may not hold in practical zero-shot transfer scenarios with large domain shifts.
@@ 4.1 Major Barriers in Prior TL Methods
As discussed, most existing TL approaches optimize a proximal objective function $f(\pi)$ and deploy its optimal policy in the target domain. However, there are limited guarantees regarding the relationship between the proxy $f(\pi)$ and the target domain value function $V^{\pi}_{P_{0}}$, except under some strong assumptions (e.g., (Chen et al., 2021)). Consequently, it is unclear whether the optimal policy for $f(\pi)$ will also perform well under $P_{0}$, leaving uncertainty in its performance. Furthermore, when multiple source domains are available, existing methods struggle to distinguish domains that are more dissimilar to the target, often treating all domains equally. This inability to prioritize relevant sources can lead to negative transfer, where the knowledge transferred from additional source domains hurt the performance of the target domain instead of improving it.
@@ 4.3 A Robust RL Based Pessimism Principle
Robust RL is inherently pessimistic when tackling environment uncertainty. Specifically, when the environment of interest lies within the defined uncertainty set, the robust value function naturally provides a conservative lower bound on performance. This property inspires our studies of robust RL in transfer learning, as constructing an uncertainty set that includes the target domain is not challenging, and thus we can obtain a conservative proxy. Generally, the learner has some prior knowledge about the domain similarities to motivate transfer learning, and such similarities are usually captured in terms of the source and target domain kernels. For instance, a common characterization is the distance/divergence between the source and target kernels, e.g., (Qu et al., 2024). Based on this information, an uncertainty set can be constructed to include the target domain kernel, and result in its robust value function as a conservative proxy.
@@
It can be shown that all of these proxies are conservative estimations of $V^{\pi}_{P_{0}}$, however, none of these proxies can be efficiently solved in a distributed setting. Specifically, (robust) DR is generally intractable and is approximated by a proximal DR objective (Jin et al., 2022). Meanwhile, solving other proxies may require sharing local source domain models or data, which is often impractical or undesirable in a distributed setting. Therefore, we propose two conservative proxies based on robust RL, and show that they can be efficiently optimized under the distributed setting.
@@ 5.1 Averaged Operator Based proxy
Notice that according to the proof of Proposition 5.4, the result holds provided that $\sigma_{\mathcal{P}_{s}^{a}}-\frac{1}{K}\sum_{k=1}^{K}\sigma_{(\mathcal{P})_{s}^{a}}(V)\leq 0$. In Addition, for $l_{p}$-norm, the support function has a duality: $\sigma(V)=\max_{\alpha}\{PV_{\alpha}+f(\Gamma,V_{\alpha})\}$ for some function $f$ (Clavier et al., 2024). Thus $\sigma_{\mathcal{P}_{s}^{a}}-\frac{1}{K}\sum_{k=1}^{K}\sigma_{(\mathcal{P})_{s}^{a}}(V)=\max_{\alpha}\{\bar{P}V_{\alpha}+f(\Gamma,V_{\alpha})\}-\frac{1}{K}\sum_{k=1}^{K}\max_{\alpha}\{P_{k}V_{\alpha}+f(\Gamma,V_{\alpha})\}\leq 0$, and the similar result still holds.
@@ 5.2 Efficient and Distributed Algorithm Design
Recall $\mathbf{T}_{k}$ being the optimal robust Bellman operator for the $k$-th source domain as Equation 4. Define the averaged optimal Bellman operator as
@@
In practice, the exact source domains may be unknown, but extensive samples from them are available, as the source domains are more free to explore. We hence assume that each local source domain can obtain an unbiased estimate of its robust Bellman operator: for any vector $Q$, an unbiased estimate $\hat{\mathbf{T}}_{k}Q$ is available to each local agent. This is a mild assumption, as there are different ways to construct an unbiased estimator of the robust Bellman operator, e.g., (Wang et al., 2023b; Kumar et al., 2023b; Yang et al., 2023; Liu et al., 2022). We also note that due to the linear structure of $\mathbf{T}_{\rm AO}$, its unbiased estimation can be directly constructed as the average of all local estimations as $\hat{\mathbf{T}}_{\rm AO}=\frac{\sum^{K}_{k=1}\hat{\mathbf{T}}_{k}}{K}.$
On the other hand, to further avoid extensive communication, the frequency of aggregation can be reduced. Specifically, we update the local estimation of each source domain individually for $E$ steps, and then aggregate all the domains once. Such a learning scheme is also used in federated or distributed learning, e.g., (Jin et al., 2022; Wang et al., 2024b; Woo et al., 2023).
@@
As the result shows, our algorithm enjoys a partial linear speedup, implying the efficiency of our algorithm for multi-domain transfer. The second term measures the trade-off between the convergence rate and communication cost, which is also observed in federated learning (FL) with heterogeneous environments (Wang et al., 2024b). We note that the convergence rate of our stochastic algorithm still enjoys a partial linear speedup, which matches the minimax optimal robust up to polylog factors (Vershynin, 2018). It hence illustrates the efficiency of our algorithm, in terms of convergence rate and communication cost.
@@ 6 Avoiding Negative Transfer: the Minimal Pessimism Principle
While our averaged operator-based method provides a conservative yet effective proxy for target domain performance, it remains vulnerable to a key challenge in multi-domain transfer: negative transfer. Similar to existing TL methods, the averaged operator-based proxy treats all source domains equally and cannot identify or down-weight the less relevant ones. Inspired by recent advances in personalized FL (Arivazhagan et al., 2019; Fallah et al., 2020; Deng et al., 2020; Tan et al., 2022), the global learner should assign higher weights to source domains that are more similar to the target domain. However, identifying these similar domains without additional information remains challenging.
@@
(1) Both ${\mathbf{T}}^{\pi}_{{\rm MP}}$ and ${\mathbf{T}}_{{\rm MP}}$ are $\gamma$-contractions, and have unique fixed points, $Q^{\pi}_{{\rm MP}}$ and $Q_{{\rm MP}}$, respectively.
(2) For any policy $\pi$, let $V^{\pi}_{{\rm MP}}(s)=\sum_{a\in\mathcal{A}}\pi(a|s)Q^{\pi}_{{\rm MP}}(s,a)$, then $V^{\pi}_{\mathcal{P}_{k}}\leq V^{\pi}_{{\rm MP}}\leq V^{\pi}_{P_{0}}$. Moreover, $V^{\pi}_{\rm AO}\leq V^{\pi}_{{\rm MP}}$.
(3) Let $\pi_{\ast}(s)=\arg\max_{a\in\mathcal{A}}Q_{{\rm MP}}(s,a)$, then $Q_{{\rm MP}}=Q^{\pi_{\ast}}_{{\rm MP}}$ and $\pi_{\ast}=\arg\max_{\pi}Q_{{\rm MP}}^{\pi}$.
@@
To address this issue, it is essential to design an alternative maximal aggregation scheme that ensures convergence despite the presence of estimation errors. We propose an aggregation strategy to overcome this challenge. Specifically, we adopt the multi-level Monte-Carlo method (Blanchet & Glynn, 2015; Blanchet et al., 2019; Wang & Wang, 2022) to construct an operator $\hat{\mathbf{M}}_{\rm MLMC}$ of the maximal aggregation based on estimated local operators. We defer the detailed construction to Section E.1.
@@
The MLMC module for MDTL-Max will increase the sample/computation complexity, which can be viewed as the price of improving effectiveness. Nevertheless, it can be reduced along two potential directions: One is to control the level number through techniques like threshold-MLMC (Wang et al., 2024f). Although it results in a biased estimation, the bias can be controlled and hence still implies convergence (see Section B.4). Another one is to relax the uncertainty set constraint. For example, for total variation, the relaxation results in the solution $P_{0}V-\Gamma\text{Span}(V)\leq\sigma(V)$ (Kumar et al., 2023a), which is much easier to compute and remains conservative. Reducing the aggregation frequency is an alternative way to reduce the computational cost, but also introduces a trade-off in the convergence rate, as we will show in the following results.